PLEASE ANSWER!!

A medicinal chemist is analyzing the interaction between a synthetic hormone (the drug norethindrone) and a hormone receptor in an effort to estimate the Kd of the hormone binding to the receptor. In addition, the chemist needs to determine the number of hormone binding sites on each receptor molecule. A series of experiments are performed in which increasing total concentrations of hormone are mixed with a fix total receptor concentration of 450nM. After binding equilibrium is reached in each experiment, the values of [L] and [RL] are determined by equilibrium dialysis.

Given the data in the table below, sketch a Scatchard plot (labeling both axes) and estimate the following parameters: Kd, Bmax, and number or binding sites on each protein (round to nearest integer).

Kd=

Bmax=

Binding sites per protein=

[L](nM) [RL] (nM)
0.06 1
9.7 150
21.8 300
35.8 450
62.3 600
112.1 800
222.9 1000
Answer #1

Scatchard equation can be written as follows.

[RL]/[L] = (-1/Kd)[RL] + Bmax/Kd ---- Equation 1

The Scatchard plot of [RL]/[L] versus [RL] can be drawn as follows.

Scatchard equation: [RL]/[L] = -0.01252*[RL] + 17.29323 ---- Equation 2

By comapring Equation 2 with Equation 1, you will get

1/Kd = 0.01252

Therefore, Kd = 79.872

The y-intercept = Bmax/Kd = 17.29323

i.e. Bmax*0.01252 = 17.29323

Therefore, Bmax = 1381.248 nm

Finally, Binding sites per protein = Bmax/Kd = 17.29323 ~ 17